A balloon starting from the ground has been ascending vertically at a uniform velocity for 4 seconds and a stone let fall from it reaches the ground in 6 seconds find the velocity of the balloon and its height when the stone was let Fall take g is equal to 10 m per second sq.

Grade:11

1 Answers

Shail patel
5 years ago
The distance  travelled by balloon is = 1/2gt2.
T=6 sec , h=1/2 ×10 × (6)2=180 mm. 
Now let the velocity of balloon be v ,
We know h=v2/2g ,
180 = v2/20 , v2 = 3600,
V = 60 m/s .
 
 Ans = height  of balloon  = 180m, 
         And velocity  is 60m/s.